# EgoEvGesture: Gesture Recognition Based on Egocentric Event Camera

### v1

@@ 5.5 Quantitatively Result on DVS128 Gesture
The competitive performance on DVS128 Gesture can be attributed to several factors inherent in our proposed architecture. First, our lightweight CNN with asymmetric depthwise convolutions effectively reduces parameters while preserving spatiotemporal features, which is crucial for handling the sparse and asynchronous nature of event-based data. This design choice not only enhances computational efficiency but also maintains the essential information needed for accurate gesture recognition. Second, the plug-and-play state-space model, acting as a context block, significantly enhances the model’s ability to filter out irrelevant motion patterns and helps focus on the actual gesture dynamics, thereby improving recognition accuracy. Third, the parameter-free Bins-Temporal Shift Module (BSTM) efficiently fuses sparse events by shifting features along bins and temporal dimensions. This mechanism is particularly beneficial for event-based data where information is spread sparsely across time and space. By effectively aggregating these sparse events, BSTM ensures that the model can capture the complete gesture patterns even from limited and dispersed event inputs. In summary, the remarkable accuracy of 96.97% on the DVS128 Gesture dataset demonstrates the strong generalization capability of our proposed architecture. This result, comparable to the state-of-the-art S-tllr approach, highlights the effectiveness of our method in handling the unique challenges of event-based data. The detailed analysis of our architecture’s components provides insights into their contributions to the overall performance, further validating the robustness and competitiveness of our approach across diverse datasets.
6 Conclusion
In this work, we address the new problem, i.e., event-based egocentric gesture recognition based on egocentric event cameras, and we introduce all the necessary tools required to address fundamental challenges in designing the method. We propose an egocentric gesture recognition framework featuring a three-stage processing pipeline: First, raw event streams are converted into Locally Normalized Event Surfaces (LNES) to suppress ego-motion artifacts from head movements. Subsequently, a lightweight Blaze encoder-decoder combined with a VMamba module extracts spatiotemporal features, addressing event sparsity and long-term temporal modeling. Finally, a parameter-free BTSM achieves motion-invariant temporal alignment through bidirectional propagation. To address the data gap, we construct EgoEvGesture, the first large-scale egocentric event-based gesture dataset, containing 5,419 samples of 38 gesture classes (single/two-hand interactions) from 10 subjects under normal and low light conditions. Experiments demonstrate that our method achieves 62.7% accuracy in heterogeneous testing with only 7M parameters, outperforming state-of-the-art approaches by 3.1%, while maintaining stable performance. As this is the first work on the egocentric gesture recognition using event cameras, the open-source release of the dataset and code is expected to foster synergistic advancements in event vision and egocentric interaction research. In the future, we plan to enhance HMS to make it more lightweight. We will leverage depth cameras as ground truth to explore egocentric hand keypoint estimation based on event cameras, and then extend the relevant findings to outdoor environments.

### v2

@@ V-E Quantitatively Result on DVS128 Gesture
Our leading recognition performance on the DVS128 Gesture dataset stems from three key architectural innovations: (1) A lightweight CNN with asymmetric depthwise convolutions that maintains spatiotemporal features while reducing parameters, crucial for sparse event-based data; (2) A plug-and-play state-space context block that focuses on gesture dynamics; (3) The parameter-free BSTM module that effectively fuses sparse spatiotemporal events. Achieving 97.0% accuracy (matching S-tllr’s reimplemented performance) validates our architecture’s generalization capability and robustness for event-based gesture recognition.
VI Conclusion
We propose a novel solution for event-based egocentric gesture recognition, addressing its key challenges through three technical contributions. In this work, we address the new problem, i.e., event-based egocentric gesture recognition based on egocentric event cameras by proposing an egocentric gesture recognition framework featuring a three-stage processing pipeline: First, raw event streams are converted into Locally Normalized Event Surfaces (LNES) to suppress ego-motion artifacts from head movements. Subsequently, a lightweight Blaze encoder-decoder combined with a VMamba module extracts spatiotemporal features, addressing event sparsity and long-term temporal modeling. Finally, a parameter-free BTSM achieves motion-invariant temporal alignment through bidirectional propagation. To address the data gap, we construct EgoEvGesture, the first large-scale egocentric event-based gesture dataset, containing 5,419 samples of 38 gesture classes (single/two-hand interactions) from 10 subjects under normal and low light conditions. Experiments demonstrate that our method achieves 62.7% accuracy in heterogeneous testing with only 7M parameters, outperforming state-of-the-art approaches by 3.1%, while maintaining stable performance. As this is the first work on egocentric gesture recognition using event cameras, the open-source release of the dataset and code is expected to foster synergistic advancements in event vision and egocentric interaction research.
In the future, we intend to enhance the head-mounted capture system to make it more lightweight. We plan to leverage depth cameras as ground truth to explore egocentric hand keypoint estimation based on event cameras, and then extend the relevant findings to outdoor environments.